OpenStudy (anonymous):

Part 1: Create your own radical expression. (1 point) Part 2: Convert it to a rational exponent. (2 points) Part 3: Explain, in complete sentences, how the expression was converted. (3 points)

hero (hero):

1. \[ \sqrt[3]{27}\] 2. \[27^{1/3}\] 3. Converted by rules of exponents and radicals

hero (hero):

\[\sqrt[a]{b} = b^{1/a}\]

OpenStudy (across):

Extending Hero's answer :P \[\sqrt[a]{x^b}=b^{b/a}\]

hero (hero):

typo across, but none the matter